OpenStudy (anonymous):

When solving the system by substitution, what would you plug into the second equation for the letter x? x+2y=4 2x+2y=6 x = 2 x = 4-2y x = 4 x = 2y+4 Solve the following system of equations by graphing. x + 2y = 4 2x + 4y = 8 What is the solution?

OpenStudy (solomonzelman):

\(\large\color{black}{ x+2y=4 \\ 2x+2y=6 }\)

OpenStudy (solomonzelman):

when you use substitution, rearrange the first equation \(\large\color{black}{ x+2y=4 }\) subtract 2y from both sides, \(\large\color{black}{ x+2y\color{red}{-2y}=4\color{red}{-2y} }\)

OpenStudy (solomonzelman):

so what will you get after subtracting 2y from both sides?

OpenStudy (anonymous):

\[x=4 ?\]

OpenStudy (anonymous):

Can you help me with another question ?

OpenStudy (solomonzelman):

you are close actually, I kept getting disconnected so couldn't reply right away

OpenStudy (solomonzelman):

\(\large\color{black}{ x+2y\color{red}{-2y}=4\color{red}{-2y} }\) \(\large\color{black}{ x\cancel{~+2y~} \cancel{~\color{red}{-2y}~}=4\color{red}{-2y} }\) \(\large\color{black}{ x=4-2y }\)

OpenStudy (solomonzelman):

so your answer for question 1 is?

OpenStudy (anonymous):

Ohh I get it now (: Thanks. And I'm not sure how to do this problem .. Solve the following system of equations by graphing. x + 2y = 8 x + 2y = -4 What is the solution

OpenStudy (solomonzelman):

you have to graph the equations and tell me where they intersect. https://www.desmos.com/calculator/e6rrarfcqe

OpenStudy (anonymous):

It looks like they are parallel ?

OpenStudy (anonymous):

Would that mean no solution ?

OpenStudy (solomonzelman):

yesm, correct. They are parallel, and thus have no solution.

OpenStudy (anonymous):

Cool! You're amazing. Thank youuu

OpenStudy (solomonzelman):

I'm not.... just helping. You are welcome!